# Train-before-Test Harmonizes Language Model Rankings

### v1

@@ Abstract
Existing language model benchmarks provide contradictory model rankings, even for benchmarks that aim to capture similar skills. This dilemma of conflicting rankings hampers model selection, clouds model comparisons, and adds confusion to a growing ecosystem of competing models. Recent work attributed ranking disagreement to the phenomenon of training on the test task: As released, different models exhibit a different level of preparation for any given test task. A candidate solution to the problem is train-before-test: Give each model the same benchmark-specific finetuning before evaluation. Our primary contribution is a broad empirical evaluation of train-before-test across 24 benchmarks and 61 models. We show that train-before-test significantly improves ranking agreement consistently across all benchmarks. Whereas rankings have little external validity to start with, they enjoy a significant degree of external validity when applying train-before-test: Model rankings transfer gracefully from one benchmark to the other. Even within the same model family, train-before-test reduces strong ranking disagreement to near-perfect agreement. In addition, train-before-test reduces the model-score matrix to essentially rank one, revealing new insights into the latent factors of benchmark performance. Our work supports the recommendation to make train-before-test a default component of LLM benchmarking [^2] Code is available at https://github.com/socialfoundations/lm-harmony..
@@ 1 Introduction
A single large language model can reliably solve many tasks. This impressive capability, however, has led to a crisis in benchmarking. A rapidly growing flurry of benchmarks tries to capture the numerous different downstream tasks that we might use a language model for. Unfortunately, different benchmarks generally give contradictory model rankings, making it difficult to compare and choose models.
Traditionally, benchmarking compared models trained for the same task on the same training data. In contrast, most large language models train on proprietary data mixes that differ from one model to the other. Recent work showed that this leads to the problem of training on the test task: the extent to which a model has encountered data similar to the test task during training confounds model comparisons, rankings, and scaling laws kaplan2020scaling . Put simply, an otherwise worse model may have simply prepared better for a specific task.
A candidate solution to the problem is to fine-tune each model on the same task-specific training data before evaluation. We call this form of evaluation train-before-test. We draw a distinction between direct evaluation and train-before-test. In direct evaluation, we evaluate the model as it ships out of the box. Under train-before-test, we fine-tune the model before evaluation. The point of fine-tuning is not to make the model better, but to level the playing field for evaluation. The hope is that train-before-test might level out task-specific model differences so as to reveal more robust model comparisons. Our primary contribution is a broad evaluation of train-before-test, which confirms that this is indeed the case.
@@ 1.1 Our Contributions
We demonstrate that the prevalent direct evaluation scheme results in strong disagreement between model ranking across various benchmarks. We show that this strong ranking disagreement persists even when restricting to benchmarks that aim to capture similar tasks. Moreover, rankings still strongly disagree when evaluating models from the same family. The situation presents a serious conundrum for model selection: Under direct evaluation, benchmarks fail to give reliable and actionable insights about which model to choose from a set of available alternatives.
@@
Train-before-test leads to ranking agreement.
@@
We broadly evaluate train-before-test on 24 benchmark datasets and 61 large language models. Train-before-test will fine-tune each model on the same task-relevant data before evaluation. Our results demonstrate that train-before-test drastically increases model ranking agreement across the board: Ranking agreement between benchmarks, measured by Kendall’s tau, improves for $95\%$ of the $\binom{24}{2}=276$ benchmark pairs. The average value of Kendall’s $\tau$ improves from $0.51$ to $0.74$. Figure 1 illustrates the result for one typical pair of benchmarks.
@@
Train-before-test aligns perplexity rankings with downstream tasks.
@@
Perplexity benchmarks used to be popular, but fell out of fashion because of the apparent disconnect between perplexity and downstream task performance (weiemergent, ; ganguli2022predictability, ; liu2023same, ). We indeed validate this disconnect when comparing model families under direct evaluation. In contrast, we demonstrate that under train-before-test evaluation, perplexity rankings align well with downstream task rankings. This creates an appealing consistency between model training objectives and downstream performance.
@@
Train-before-test sheds light on the latent factors of benchmark performance.
@@
Consider the large benchmark-model matrix, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$. Several works have considered this matrix for different reasons and found that it is approximately low rank (Ruan2024ObservationalSL, ; Owen2024HowPI, ; Burnell2023RevealingTS, ), but not quite. The first singular value is dominant and correlates with pre-training compute. However, the other components aren’t negligible and it wasn’t clear how to interpret them.
We show that train-before-test clarifies this state of affairs. After train-before-test, the benchmark-model matrix is essentially rank $1$. The first principal component accounts for $85\%$ of the explained variance across all models, and for $93\%$ of the variance for a single model family. This leads us to speculate that the lower components of the direct evaluation matrix correspond to how much a model was trained on a given kind of task. This would be a compelling reason why the lower singular values shrink under train-before-test.
@@ 2 Related Work
Benchmarking has played a central role in the advancement of machine learning (liberman2010obituary, ; hardt2022patterns, ). While absolute model performance is often fragile to even seemingly minor changes in evaluation data (candela2009dataset, ; torralba2011unbiased, ; albadawy2018deep, ; taori2020measuring, ; Tsipras2020FromIT, ; miller2020effect, ), relative model performance—that is, model rankings—tends to transfer surprisingly well across classical benchmarks (Yadav2019ColdCT, ; recht2019imagenetclassifiersgeneralizeimagenet, ; miller2020effect, ). For instance, prior work (kornblith2019better, ; barbu2019objectnet, ) has shown that model rankings on ImageNet (deng2009imagenet, ) also transfer to other image classification and object recognition benchmarks. Moreover, Salaudeen and Hardt (2024, (salaudeen2024imagenot, )) demonstrated that ImageNet rankings remain robust even under major dataset variations (salaudeen2024imagenot, ). This transferability of model rankings is highly desirable, as it indicates that progress on specific benchmarks reliably reflects broader scientific advancements (liao2021we, ; hardt2025emerging, ).
@@
However, the emergence of foundation models has dramatically transformed the benchmarking landscape compared to the ImageNet era (liang2022holistic, ; srivastava2022beyond, ; Weidinger2025TowardAE, ). With huge training costs and much improved capabilities (qwen, ; gemma, ; llama, ; Ramesh2021ZeroShotTG, ; Anil2023GeminiAF, ; openai2023GPT4, ), practitioners now lean towards directly evaluating LLMs across a wide range of different benchmarks, in the hope of obtaining a more comprehensive assessment of their capabilities (liang2022holistic, ; bbh, ; mmlu, ; open-llm-leaderboard, ; open-llm-leaderboard-v2, ). This shift introduces new challenges, as model rankings across different tasks may vary significantly (liang2022holistic, ; huan2025does, ; lourie2025scaling, ). Zhang and Hardt (2024, (zhang2024inherent, )) draw an analogy between multi-task benchmarks and voting systems (arrow2012social, ), revealing that a multi-benchmarking approach with diverse rankings inherently lacks robustness to minor changes and thus cannot provide a stable unified ranking.
@@
This lack of unified ranking is sometimes seen as a desirable feature within the community (liang2022holistic, ). Some argue that variability reflects the multifaceted strengths and weaknesses of LLMs, suggesting that users should select the best model tailored to their specific needs (ghosh2024onebench, ; Zhang2024TaskMA, ; Shnitzer2023LargeLM, ). For example, a user who focuses on mathematical tasks could prioritize the math benchmark to choose the optimum model. However, there are two significant concerns regarding this approach: First, the user-driven selection strategy poses challenges for model developers. Given the resource-intensive nature of LLM development (guo2025deepseek, ), it is impractical to release a different model for every potential use case. Moreover, developers typically aim to create a general-purpose model (qwen, ; llama, ; gemma, ); however, such a desideratum is often difficult to reliably measure due to the inconsistent rankings observed across benchmarks. Second, we demonstrate in this paper that benchmarks within the same task category can still exhibit substantial discrepancies in model rankings.
@@
One potential reason for the observed inconsistencies in model rankings is that models vary substantially in their training data (Gadre2023DataCompIS, ; albalaksurvey, ). In particular, Dominguez et al. (2024, (dominguez2024training, )) show that models vary in their degree of preparedness for popular benchmarks. Building on this idea, we introduce the notion of train-before-test, wherein we fine-tune each model on the corresponding training set to get every model well-prepared. We then investigate how this approach improves the consistency of rankings across benchmarks and discuss its implications for future benchmarking practices.
@@ 3 Experiments
In this section, we examine the cross-benchmark ranking agreement of 61 language models across 24 benchmarks. We find that ranking agreement tends to be low, with an average Kendall’s $\tau$ of 0.51. We then examine benchmark agreement under a different benchmarking methodology, which we refer to as train-before-test. Specifically, we fine-tune on a benchmark’s train set prior to evaluating on said benchmark. Compared to direct evaluation, train-before-test improves cross-benchmark ranking agreement on 95% of the benchmark pairs considered. The improvements in ranking agreement are typically large, with the average Kendall’s $\tau$ increasing to 0.74.
We additionally find that train-before-test greatly improves the agreement between perplexity rankings and downstream benchmarks. This result holds consistently across three general domain corpus, newly collected from Wiki, arXiv, and StackExchange. The average Kendall’s $\tau$ improves from 0.47 to 0.73 with train-before-test, leading to much better consistency between the training objective and downstream benchmark performance.
Finally, we discuss the implications of increased cross-benchmark agreement. In doing so, we perform Principal Component Analysis (PCA) over the model score matrix comprising all benchmark scores. We then analyze its principal components both under direct model evaluation and train-before-test. We find that train-before-test greatly increases the share of variance explained by the first principal component (PC1), from 69% for direct evaluation to 85% for train-before-test. In both cases, PC1 aligns well with model pre-training compute. By further controlling for model family and only considering Qwen models, we show that the explained variance ratio of PC1 further increases to 93%, making the model-score matrix essentially rank one.
@@
Figure 2: Mean ranking agreement between each benchmark and all others. We calculate Kendall’s $\tau$ between each benchmark and every other benchmark, and then average the results. Compared to direct evaluation, train-before-test consistently improves ranking agreement, often by a large margin. A detailed comparison of Kendall’s $\tau$ values for every benchmark pair is provided in Appendix B.1. On average, the overall average Kendall’s $\tau$ is 0.51 for direct evaluation and 0.74 for train-before-test.
@@
Figure 3: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right). We categorize benchmarks into language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med), see Table 1. Kendall’s $\tau$ is averaged across all pairs of benchmarks that belong to two given categories. The diagonal entries represent intra-category agreement and the other entries represent inter-category agreement. Train-before-test improves both intra- and inter-category ranking agreement in all instances.
3.1 Experiment Setting
Benchmark selection.
Generation-based benchmarks are often computationally intensive to evaluate, as base models typically generate text until reaching their maximum sequence length. We select only NQ-Open and GSM8K from the generation-based benchmarks, excluding the remaining seven due to computational constraints. Among the likelihood-based benchmarks, we further exclude six due to observed anomalies during fine-tuning, such as a lack of performance improvement in over 20% of models. See Appendix A.1 for additional details.
Our final selection consists of 24 benchmarks covering diverse domains and task types. These benchmarks are primarily multiple-choice question answering benchmarks, with accuracy as the task metric. We categorize all benchmarks by their descriptions, see Table 1.
If a benchmark does not come with a validation split, we randomly allocate 20% of the training data as the validation set. To save computational resources, we cap the number of training data at 50,000, validation data at 1,000, and testing data at 10,000.
Model selection.
We consider 61 language models across six model families: Llama (llama, ), Qwen (qwen, ), Gemma (gemma, ), Pythia (pythia, ), GPT-2 (gpt2, ) and Yi (yi, ). Due to computational constraints, we select models with fewer than 14B parameters. See Table 2 for the full list. We include both base and instruction-tuned models, and use the suffix -IT to denote instruction-tuned models.
Evaluation setup.
We evaluate the 61 models across all 24 benchmarks using both direct evaluation and train-before-test evaluation. We use the lm-eval-harness library for evaluation. We evaluate models zero-shot (Brown2020LanguageMA, ). For direct evaluation, we simply evaluate the model as it is. For train-before-test, we fine-tune models for five epochs, and select the best performing checkpoint based on performance on a separate validation set. We therefore fine-tune a total of $61$ models on $24$ benchmarks, totaling $61\times 24=1,464$ fine-tuned models. We use parameter-efficient fine-tuning (PEFT) (Hu2021LoRALA, ; peft, ), with fixed hyperparameters. See more details in Appendix A.3. Each fine-tuned model is then evaluated on the corresponding benchmark’s test set. For each benchmark, we rank models according to their performance. We then measure the ranking correlation across pairs of benchmarks using Kendall’s $\tau$ (kendall1938new, ).

### v2

@@ Abstract
Existing language model benchmarks provide contradictory model rankings, even for benchmarks that aim to capture similar skills. This dilemma of conflicting rankings hampers model selection, clouds model comparisons, and adds confusion to a growing ecosystem of competing models. In this paper, we take a different perspective on model comparison: instead of relying on out-of-the-box performance via direct evaluation, we compare model potential by providing each model with identical benchmark-specific fine-tuning before evaluation. We call this approach train-before-test. Our primary contribution is a comprehensive empirical evaluation of model potential across 24 benchmarks and 61 models. First, we demonstrate that model potential rankings obtained through train-before-test exhibit remarkable consistency across all benchmarks. Whereas traditional rankings demonstrate little external validity under direct evaluation, they enjoy a significant degree of external validity when applying train-before-test: model potential rankings transfer gracefully from one benchmark to another. Second, train-before-test restores the connection between perplexity and downstream task performance, lost under direct evaluation. Remarkably, even pre-finetuning perplexity of a base model predicts post-finetuning downstream performance, suggesting that ranking consistency reflects inherent model potential rather than fine-tuning artifacts. Finally, train-before-test reduces the model-score matrix to essentially rank one, indicating that model potential is dominated by one latent factor, uncovered by train-before-test. Our work supports the recommendation to make train-before-test a default component of LLM benchmarking [^2] Code is available at https://github.com/socialfoundations/lm-harmony..
@@ 1 Introduction
Existing language model benchmarks provide contradictory model rankings, even for benchmarks that aim to capture similar skills (liang2022holistic, ; open-llm-leaderboard, ; open-llm-leaderboard-v2, ). This inconsistency poses a serious challenge: how can we reliably compare, rank, and select models when different benchmarks yield conflicting information? While this ranking disagreement is often attributed to the diverse capabilities of large language models (Ruan2024ObservationalSL, ), it creates a conundrum in practice that muddles model development decisions (zhang2024inherent, ).
Current evaluation methodology works from direct evaluation, probing models via black-box function calls. However, large language models are trained on diverse, often proprietary data mixes that vary significantly across models (llama, ; gemma, ; Guha2025OpenThoughtsDR, ). Recent work showed that this leads to the problem of training on the test task (dominguez2024training, ): the extent to which a model has encountered data similar to the test task during training confounds model comparisons, rankings, and scaling laws kaplan2020scaling . Put simply, an otherwise inferior model may have simply prepared better for a specific task.
In this paper, we take a fresh perspective on evaluation methodology: in contrast with direct evaluation, we compare model potential by giving each model the same task-specific fine-tuning. We call this approach train-before-test. Its goal is to achieve valid model comparisons by ensuring that all models receive equal preparation for the test.
We envision train-before-test as a tool for regret-free model selection for downstream applications. Increasingly, practitioners select one from many available models with the goal of adapting for a specific task. Under direct evaluation the best model to begin with may no longer be the best model after task-specific preparation. In contrast, we show that train-before-task yields model comparisons and rankings that enjoy broad external validity.
@@ 1.1 Our Contributions
We demonstrate that the prevalent direct evaluation scheme results in strong disagreement between model ranking across various benchmarks. We show that this strong ranking disagreement persists even when restricting to benchmarks that aim to capture similar tasks. Moreover, rankings still strongly disagree when evaluating models from the same family. The situation presents a serious conundrum for model selection: Under direct evaluation, benchmarks fail to give reliable and actionable insights for model choosing among multiple alternatives.
@@
Train-before-test leads to consistent model potential rankings.
@@
We comprehensively evaluate train-before-test across 24 benchmark datasets and 61 large language models. By fine-tuning each model on identical task-relevant data before evaluation, we uncover remarkably consistent model potential rankings. Ranking agreement between benchmarks, measured by Kendall’s tau, improves for 274 out of 276 benchmark pairs, with the average Kendall’s $\tau$ increasing from 0.52 to 0.76. Figure 1 illustrates the result for one typical pair of benchmarks. This consistency suggests that model potential, unlike out-of-the-box performance, has external validity (Salaudeen2025MeasurementTM, ) and transfers gracefully across different tasks.
@@
Model potential aligns perplexity rankings with downstream tasks.
@@
Perplexity benchmarks used to be popular, but fell out of fashion because of the apparent disconnect between perplexity and downstream task performance (weiemergent, ; ganguli2022predictability, ; liu2023same, ; Magnusson2023PalomaAB, ; Lourie2025ScalingLA, ). We indeed validate this disconnect when comparing model families under direct evaluation. However, train-before-test restores this fundamental relationship in two ways. First, we show that post-fine-tuning perplexity rankings align well with post-fine-tuning downstream task rankings, creating consistency between training objectives and task performance. Second, and more remarkably, for base (non-instruction-tuned) models, even pre-fine-tuning perplexity predicts post-fine-tuning downstream performance. This suggests that the ranking consistency we observe reflects inherent model potential rather than artifacts of fine-tuning.
@@
Train-before-test sheds light on the latent factors of benchmark scores.
@@
Consider the large benchmark-model score matrix, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$. Several works have considered this matrix for different reasons and found that it is approximately low rank (Ruan2024ObservationalSL, ; Owen2024HowPI, ; Burnell2023RevealingTS, ), but not quite. The first singular value is dominant and correlates with pre-training compute. However, the other components aren’t negligible, and their interpretation remains unclear. We show that train-before-test clarifies this state of affairs. After train-before-test, the benchmark-model matrix is essentially rank one. The first principal component accounts for $86\%$ of the explained variance across all models, and for $93\%$ of the variance for a single model family. This suggests that model potential is dominated by a single latent factor, while the additional components observed in direct evaluation may reflect task-specific training exposure.
@@ 2 Related Work
Benchmarking has played a central role in the advancement of machine learning (liberman2010obituary, ; hardt2022patterns, ). While absolute model performance is often fragile to even seemingly minor changes in evaluation data (candela2009dataset, ; torralba2011unbiased, ; albadawy2018deep, ; taori2020measuring, ; Tsipras2020FromIT, ), relative model performance—that is, model rankings—tends to transfer surprisingly well across classical benchmarks (Yadav2019ColdCT, ; recht2019imagenetclassifiersgeneralizeimagenet, ; miller2020effect, ). For instance, prior work (kornblith2019better, ; barbu2019objectnet, ) has shown that model rankings on ImageNet (deng2009imagenet, ) also transfer to other image classification and object recognition benchmarks. Moreover, Salaudeen and Hardt (2024, (salaudeen2024imagenot, )) demonstrated that ImageNet rankings remain robust even under major dataset variations. This transferability of model rankings is highly desirable, as it indicates that progress on specific benchmarks reliably reflects broader scientific advancements (liao2021we, ; hardt2025emerging, ).
@@
However, the emergence of foundation models has dramatically transformed the benchmarking landscape compared to the ImageNet era (liang2022holistic, ; srivastava2022beyond, ; Weidinger2025TowardAE, ). With huge training costs and much improved capabilities (qwen, ; llama, ; Ramesh2021ZeroShotTG, ; Anil2023GeminiAF, ; openai2023GPT4, ), practitioners now lean towards directly evaluating LLMs across a wide range of different benchmarks, in the hope of obtaining a more comprehensive assessment of their capabilities (liang2022holistic, ; bbh, ; mmlu, ; open-llm-leaderboard, ; open-llm-leaderboard-v2, ). This shift introduces new challenges, as model rankings across different tasks may vary significantly (huan2025does, ; lourie2025scaling, ). Zhang and Hardt (2024, (zhang2024inherent, )) draw an analogy between multi-task benchmarks and voting systems (arrow2012social, ), revealing that a multi-task benchmarking approach with diverse rankings inherently lacks robustness to minor changes and thus cannot provide a stable unified ranking.
@@
This lack of unified ranking is sometimes seen as a desirable feature within the community (liang2022holistic, ). Some argue that variability reflects the multifaceted strengths and weaknesses of LLMs, suggesting that users should select the best model tailored to their specific needs (ghosh2024onebench, ; Zhang2024TaskMA, ; Shnitzer2023LargeLM, ). For example, a user who focuses on mathematical tasks could prioritize the math benchmark to choose the optimum model. However, there are two significant concerns regarding this approach: First, the user-driven selection strategy poses challenges for model developers. Given the resource-intensive nature of LLM development (guo2025deepseek, ), it is impractical to release a different model for every potential use case. Moreover, developers typically aim to create a general-purpose model (qwen, ; llama, ); however, such a desideratum is often difficult to reliably measure due to the inconsistent rankings observed across benchmarks. Second, we demonstrate in this paper that benchmarks within the same task category can still exhibit substantial discrepancies in model rankings.
@@
One potential reason for the observed inconsistencies in model rankings is that models vary substantially in their training data (Gadre2023DataCompIS, ; albalaksurvey, ). In particular, Dominguez et al. (2024, (dominguez2024training, )) show that models vary in their degree of preparedness for popular benchmarks. Building on this idea, we introduce the notion of train-before-test, wherein we fine-tune each model on the corresponding training set so every model arrives well prepared. While extensive literature exists on investigating different fine-tuning strategies for LLMs (Lin2024SelectingLL, ; Zeng2025LENSLLMUF, ; Lester2021ThePO, ), this lies outside the scope of our investigation. Instead, we apply standardized fine-tuning (peft, ) as an evaluation tool to give all models equivalent preparation before testing. Rather than studying models from the same family with varying pre-training compute (Zhang2024WhenSM, ; kaplan2020scaling, ), our experiments cover 61 models from six families and 24 tasks from different categories. We study how train-before-test improves ranking consistency across benchmarks and its implications for benchmarking practices.
@@ 3 Experiments
In this section, we examine the cross-benchmark ranking agreement of 61 language models across 24 benchmarks. We find that ranking agreement tends to be low, with an average Kendall’s $\tau$ of 0.52. We then examine benchmark agreement under a different benchmarking methodology, which we refer to as train-before-test. Specifically, we fine-tune on a benchmark’s train set prior to evaluating on said benchmark. Compared to direct evaluation, train-before-test improves cross-benchmark ranking agreement on almost all benchmark pairs considered. The improvements in ranking agreement are typically large, with the average Kendall’s $\tau$ increasing to 0.76.
We additionally find that train-before-test greatly improves the agreement between perplexity rankings and downstream benchmarks. This result holds consistently across three general domain corpus, newly collected from Wiki, arXiv, and StackExchange. We retrained content only from 2025 to ensure models had not seen those texts during pretraining. The average Kendall’s $\tau$ between perplexity ranking and 24 downstream task rankings improves from 0.48 to 0.74 with train-before-test, leading to much better consistency between the training objective and downstream benchmark performance. Moreover, for base models, pre-fine-tuning perplexity ranking remains consistent with post-fine-tuning downstream rankings (average Kendall’s $\tau$ = 0.78). This consistency does not hold for instruction-tuned models.
Finally, we discuss the implications of increased cross-benchmark agreement. In doing so, we perform Principal Component Analysis (PCA) over the model score matrix comprising all benchmark scores. We then analyze its principal components both under direct model evaluation and train-before-test. We find that train-before-test greatly increases the share of variance explained by the first principal component (PC1), from 70% for direct evaluation to 86% for train-before-test. In both cases, PC1 aligns well with model pre-training compute. By further controlling for model family and only considering Qwen models, we show that the explained variance ratio of PC1 further increases to 93%, making the model-score matrix essentially rank one.
3.1 Experiment Setting
@@
Benchmark selection.
Generation-based benchmarks are often computationally intensive to evaluate, as base models typically generate text until reaching their maximum sequence length. These benchmarks are also over-challenging for smaller models with limited parameters, such as GPT-2 (gpt2, ). Therefore, we select only NQ-Open and GSM8K from the generation-based benchmarks. Among the likelihood-based benchmarks, we further exclude six due to observed anomalies during fine-tuning, such as a lack of performance improvement in over 20% of models. See Appendix A.1 for details.
Our final selection consists of 24 benchmarks covering diverse domains and task types. These benchmarks are primarily multiple-choice question answering benchmarks, with accuracy as the task metric. We categorize all benchmarks by their descriptions, see Table 1. If a benchmark does not come with a validation split, we randomly allocate 20% of the training data as the validation set. To save computational resources, we cap the number of training data at 50,000, validation data at 1,000, and testing data at 10,000.
Table 2: Models considered, categorized by model family.
Family Model Name Suffix
Llama-
3-8B, 3.1-8B, 3.2-1B, 3.2-3B, 3-8B-IT, 3.1-8B-IT, 3.2-1B-IT, 3.2-3B-IT
Qwen-
1.5-0.5B, 1.5-1.8B, 1.5-4B, 1.5-7B, 1.5-14B, 2-0.5B, 2-1.5B, 2-7B, 2.5-0.5B, 2.5-1.5B, 2.5-3B, 2.5-7B,
2.5-14B, 1.5-0.5B-IT, 1.5-1.8B-IT, 1.5-4B-IT, 1.5-7B-IT, 1.5-14B-IT, 2-0.5B-IT, 2-1.5B-IT, 2-7B-IT,
2.5-0.5B-IT, 2.5-1.5B-IT, 2.5-3B-IT, 2.5-7B-IT, 2.5-14B-IT
Gemma-
2B, 7B, 2-2B, 2-9B, 2B-IT, 7B-IT, 2-2B-IT, 2-9B-IT
GPT2-
124M, 335M, 774M, 1.5B
Pythia-
70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B
Yi-
6B, 9B, 6B-IT, 1.5-6B, 1.5-9B, 1.5-6B-IT, 1.5-9B-IT
Model selection.
We consider 61 language models across six model families: Llama (llama, ), Qwen (qwen, ), Gemma (gemma, ), Pythia (pythia, ), GPT-2 (gpt2, ) and Yi (yi, ). Due to computational constraints, we select models with no more than 14B parameters. See Table 2 for the full list. We include both base and instruction-tuned models, and use the suffix -IT to denote instruction-tuned models.
Evaluation setup.
We evaluate the 61 models across all 24 benchmarks using both direct evaluation and train-before-test evaluation. We use the lm-eval-harness library for evaluation. We evaluate models zero-shot (Brown2020LanguageMA, ). For direct evaluation, we simply evaluate the model as it is. For train-before-test, we fine-tune models for five epochs using learning rates in $\{1\mathrm{e}-5,2\mathrm{e}-5,5\mathrm{e}-5\}$, separately. The best performing checkpoint is then selected based on performance on a separate validation set, yielding $61\times 24=1,464$ fine-tuned models in total. We use parameter-efficient fine-tuning (PEFT) (Hu2021LoRALA, ; peft, ). See more details in Appendix A.2. Each fine-tuned model is then evaluated on the corresponding benchmark’s test set. For each benchmark, we rank models according to their performance. We then measure the ranking correlation across pairs of benchmarks using Kendall’s $\tau$ (kendall1938new, ).
Figure 2: Mean ranking agreement between each benchmark and all others. We calculate Kendall’s $\tau$ between each benchmark and every other benchmark, and then average the results. Compared to direct evaluation, train-before-test consistently improves ranking agreement, often by a large margin. A detailed comparison of Kendall’s $\tau$ values for every benchmark pair is provided in Appendix B.1. On average, the overall average Kendall’s $\tau$ is 0.52 for direct evaluation and 0.76 for train-before-test.
@@
Figure 3: Cross-category ranking agreement for direct evaluation (left) and train-before-test (right). We categorize benchmarks into language understanding (LU), commonsense reasoning (CR), question answering (QA), physics/biology/chemistry (PBC), math (Math), and medicine (Med), see Table 1. Kendall’s $\tau$ is averaged across all pairs of benchmarks that belong to two given categories. The diagonal entries represent intra-category agreement and the others represent inter-category agreement. Train-before-test improves both intra- and inter-category ranking agreement in all instances.